# Kernel Mean Embedding Topology: Weak and Strong Forms for Stochastic Kernels and Implications for Model Learning

### v1

@@ Abstract
We introduce a novel topology, called Kernel Mean Embedding Topology, for stochastic kernels, in a weak and strong form. This topology, defined on the spaces of Bochner integrable functions from a signal space to a space of probability measures endowed with a Hilbert space structure, allows for a versatile formulation. This construction allows one to obtain both a strong and weak formulation. (i) For its weak formulation, we highlight the utility on relaxed policy spaces, and investigate connections with the Young narrow topology and Borkar (or $w^{*}$)-topology, and establish equivalence properties. We report that, while both the $w^{*}$-topology and kernel mean embedding topology are relatively compact, they are not closed. Conversely, while the Young narrow topology is closed, it lacks relative compactness. (ii) We show that the strong form provides an appropriate formulation for placing topologies on spaces of models characterized by stochastic kernels with explicit robustness and learning theoretic implications on optimal stochastic control under discounted or average cost criteria. (iii) We show that this topology possesses several properties making it ideal to study optimality, approximations, robustness and continuity properties. In particular, the kernel mean embedding topology has a Hilbert space structure, which is particularly useful for approximating stochastic kernels through simulation data.
6,
@@ 1 Weak and Strong Topologies on Stochastic Kernels
$\Gamma=\bigg{\{}\gamma:\text{$\gamma$ is a measurable function from ${\mathsf{Y}}$ to ${\mathcal{P}}({\mathsf{U}})$}\bigg{\}},$
@@
Our primary area of application will be control theory: In both deterministic and stochastic control theory, such stochastic kernels play a crucial role, representing the dynamics of a stochastic system ([27, Lemma 1.2], [13, Lemma 3.1], or [3, Lemma F]) or the sets of relaxed/randomized control policies ([66],[23, 36, 35],[12]).
@@ 1.1 Summary of Main Results
Figure 1: Hierarchy of topologies on stochastic kernels and relations with the topologies introduced in our paper: The strong kernel mean embedding topology and weak kernel mean embedding topology: (a) The Young-narrow and Borkar topologies are known in the literature, and their equivalence on the set of stochastic kernels is studied in [68]. (b) In this paper, we introduce the weak kernel mean embedding topology and demonstrate its equivalence to the Young-narrow and Borkar topologies (see Theorem 3.10). (c) We also introduce a strong version of the kernel mean embedding topology, which has been used to establish empirical consistency results for approximating the conditional mean embeddings of stochastic kernels [61, 42]. It clearly implies the weak version of the kernel mean embedding topology. (d’) As established in Theorem 4.1, convergence in the weak kernel mean embedding topology does not imply convergence in the strong version. (d) However, as shown in Theorem 4.2, under the assumption of bounded and equicontinuous densities for the probability measures in the stochastic kernels, we can establish the implication from the weak to the strong version of the kernel topology. (e) Conditional mean embedding framework [58] embeds probabilities of stochastic kernels into reproducing kernel Hilbert spaces (RKHSs), treating stochastic kernels as operators between two Hilbert spaces. Initially, these spaces were assumed to be the RKHSs on the input and output spaces. However, recent work [33] shows that the input RKHS must be replaced by an $L_{2}$-space, as the conditional expectation of an output function belongs to this $L_{2}$-space, not the input RKHS. This results in an operator norm topology for stochastic kernels. [42, Theorem 3.6] shows the strong kernel mean embedding topology dominates this topology. (f) Alternatively, stochastic kernels can be viewed as Hilbert-Schmidt operators between an RKHS and an $L_{2}$-space, where the Hilbert-Schmidt topology is equivalent to the strong kernel mean embedding topology [42, Theorem 5.1].
@@ 1.1.1 A Weak Formulation
The Young Narrow Topology. Since Young’s seminal paper [66], a particularly prominent approach has involved studying the topologies on Young measures defined by stochastic kernels. In this framework, stochastic kernels are identified with probability measures defined on a product space, maintaining a fixed marginal in the input space – typically taken to be the Lebesgue measure in the context of optimal deterministic control (see [66, 39],[15, Section 2.1],[64, p. 254],[38], [4, Theorem 2.2]).
Thus, under the Young narrow topology, one associates with $\Gamma$ the set of probability measures induced on the product space ${\mathsf{Y}}\times{\mathsf{U}}$ with a fixed marginal $\mu$ on ${\mathsf{Y}}$. On this product space, several weak topologies and their equivalence properties have been studied in the literature; see e.g, [5], [4, Theorem 2.2] and [24, Theorem 3.32].
The generalization to stochastic control problems to encompass more diverse input measures has become a common practice, especially in applications involving partially observed and decentralized stochastic control. Many of the references have adopted this approach to define control topologies. For example, this methodology has been used in studies on existence and approximation results related to optimal stochastic control [23, 36, 35], decentralized stochastic control [67, Theorem 5.2] and [53, Section 4], piecewise deterministic optimal stochastic control [7, 6], economics and game theory [41, 40, 4], mean-field control policies [8], and optimal quantization [69, Definition 2.1].
The $w^{*}$ or Borkar Topology. Another notable topology on stochastic kernels is the one introduced by Borkar for randomized controls [12]. This topology, which is further elaborated upon in [2, Section 2.4] and builds on the work in [10], is formulated as a $w^{*}$-topology on stochastic kernels. In this framework, policies are treated as mappings from states or measurements to the space of signed measures with bounded variation, denoted as ${\mathcal{M}}({\mathsf{U}})$, where probability measures ${\mathcal{P}}({\mathsf{U}})$ form a subset. We also refer the reader to [19, 21] for further references on such a $w^{*}$-formulation on stochastic kernels, in particular when instead of countably additive signed measures, finitely additive such measures are also considered.
@@
Under the Borkar topology, one studies the space $\Gamma$ using a $w^{*}$-topology formulation as a bounded subset of the set of maps from ${\mathsf{Y}}$ to the space of signed measures with finite variation, which can be viewed as the topological dual of continuous functions that vanish at infinity. Therefore, according to the Banach-Alaoglu theorem [25, Theorem 5.18], this leads to a relatively sequentially compact metric space. Specifically, the unit ball of $L_{\infty}(\mu,{\mathcal{M}}({\mathsf{U}}))=\mathcal{L}_{1}(\mu,C_{0}({\mathsf{U}}))^{*}$ is compact under the $w^{*}$-topology, resulting in a relatively compact metric topology on stochastic kernels. It is important to note that the presentations in [12, Section 3] and [2, Section 2.4] differ slightly, although the induced topology remains identical. An equivalent representation of this topology is provided in [2, Lemma 2.4.1] (see also [12, Lemma 3.1]).
@@
For our weak formulation, we will obtain several structural properties and investigate its connections with the Young narrow topology and Borkar (or $w^{*}$)-topology and establish equivalence properties with both the Young narrow topology and the $w^{*}$-topology on the set of stochastic kernels. We demonstrate that the new topology we introduce is, in fact, equivalent to both the Young narrow topology and the $w^{*}$-topology on the set of stochastic kernels. This equivalence will be rigorously defined later in the paper, but the idea is that while these topologies may appear slightly different in their construction, they share fundamental properties when applied to stochastic kernels. Such a result is significant because each of these topologies offers distinct advantages in various contexts. For example, based on the results we present, both the $w^{*}$-topology and the kernel mean embedding topology are relatively (sequentially) compact, which is a crucial property when proving existence results. However, they are not closed. In contrast, the Young narrow topology is closed but lacks relative (sequential) compactness. Interestingly, to make the $w^{*}$-topology and the kernel mean embedding topology closed, a tightness condition is required – this prevents stochastic kernels from losing mass at infinity. Similarly, Prokhorov’s theorem shows that the same tightness condition is necessary for the Young narrow topology to achieve relative compactness. Thus, in terms of proving existence results via Weierstrass’s extreme value theorem, these three topologies are equally advantageous. However, the $w^{*}$-topology offers a valuable duality structure that can be leveraged for certain analyses. In addition, the kernel mean embedding topology possesses a Hilbert space structure, which can facilitate approximations of stochastic kernels via simulations, as shown in the work [42].
@@ 1.1.2 A Strong Formulation
Initially, stochastic kernels were assumed to act as operators between the RKHSs of the input and output spaces. However, recent findings [33] show that the input RKHS must be replaced by an $L_{2}$-space because the conditional expectation of an output function belongs to this $L_{2}$-space, not the input RKHS. This shift leads to an operator norm topology for stochastic kernels. Notably, the strong kernel mean embedding topology dominates the operator norm topology [42, Theorem 3.6]. Alternatively, stochastic kernels can be viewed as Hilbert-Schmidt operators between an RKHS and an $L_{2}$-space. In this context, the Hilbert-Schmidt topology is equivalent to the strong kernel mean embedding topology [42, Theorem 5.1]. While exact representations of stochastic kernels via covariance and cross-covariance operators are generally infeasible [33], certain conditions enable such representations using pseudo-inverses and adjoints of the operators [34, Theorem 5.8]. When these conditions are not met, approximate representations can be constructed using Tikhonov-Phillips regularization techniques [42, Section 7].
@@ 1.2 Notation and Conventions
[…]
@@ 2 Bochner Spaces and Their Duals
[…]
@@
We now identify the topological dual of $L_{p}\bigl{(}\mu,{\mathsf{B}}\bigr{)}$ which is denoted by $L_{p}\bigl{(}\mu,{\mathsf{B}}\bigr{)}^{*}$. We start with the definition of weakly-measurable functions from ${\mathsf{Y}}$ to ${\mathsf{B}}^{*}$. A function $\gamma:{\mathsf{Y}}\rightarrow{\mathsf{B}}^{*}$ is called weakly-measurable [16, p. 18] if the mapping ${\mathsf{Y}}\ni y\mapsto\langle\gamma(y),b\rangle\in\mathbb{R}$ is ${\mathcal{Y}}/{\mathcal{B}}(\mathbb{R})$-measurable for all $b\in{\mathsf{B}}$. Let $\mathcal{L}\bigl{(}\mu,{\mathsf{B}}^{*}\bigr{)}$ denote the set of all weakly-measurable functions. Then, we define the following subsets
[…]
@@
[16, Theorem 1.5.5] Let $q$ be the conjugate of $p$; i.e., $11/p+1/q=$ 1. For any $\gamma\in\mathcal{L}_{q}\bigl{(}\mu,{\mathsf{B}}^{*}\bigr{)}$ and $f\in L_{p}\bigl{(}\mu,{\mathsf{B}}\bigr{)}$, let
@@
If ${\mathsf{B}}^{*}$ is reflexive, then $L_{p}\bigl{(}\mu,{\mathsf{B}}\bigr{)}^{*}$ can be identified with set of $q$-Bochner integrable functions $L_{q}\bigl{(}\mu,{\mathsf{B}}^{*}\bigr{)}$ [46, Theorem 4.2.26]. Indeed, in this case, by [46, Corollary 4.2.5], we have
@@
that is weakly measurability is equivalent to strong measurability111[46, Corollary 4.2.5] states that weak measurability is equivalent to strong measurability if $B^{*}$ is separable. But if ${\mathsf{B}}^{*}$ is reflexive, then ${\mathsf{B}}$ is also reflexive. Hence $({\mathsf{B}}^{*})^{*}={\mathsf{B}}$, and so, weak measurability is equivalent to weak measurability. Since ${\mathsf{B}}$ is assumed to be separable, ${\mathsf{B}}^{*}$ is also separable. Therefore, weak-measurability is equivalent to strong measurability..
@@
[…]
If $G$ is a closed and bounded subset, then $\mathcal{L}_{q}\bigl{(}\mu,G\bigr{)}$ is also closed and bounded in $\mathcal{L}_{q}\bigl{(}\mu,{\mathsf{B}}^{*}\bigr{)}$, and so, by Banach-Alaoglu Theorem [25, Theorem 5.18], it is compact with respect to weak-topology. Since $L_{p}\bigl{(}\mu,{\mathsf{B}}\bigr{)}$ is separable, by [30, Lemma 1.3.2], $\mathcal{L}_{q}\bigl{(}\mu,G\bigr{)}$ is metrizable, and so, is also sequentially compact.

### v2

@@ Abstract
We introduce a novel topology, called Kernel Mean Embedding Topology, for stochastic kernels, in a weak and strong form. This topology, defined on the spaces of Bochner integrable functions from a signal space to a space of probability measures endowed with a Hilbert space structure, allows for a versatile formulation. This construction allows one to obtain both a strong and weak formulation. (i) For its weak formulation, we highlight the utility on relaxed policy spaces, and investigate connections with the Young narrow topology and Borkar (or $w^{*}$)-topology, and establish equivalence properties. We report that, while both the $w^{*}$-topology and kernel mean embedding topology are relatively compact, they are not closed. Conversely, while the Young narrow topology is closed, it lacks relative compactness. (ii) We show that the strong form provides an appropriate formulation for placing topologies on spaces of models characterized by stochastic kernels with explicit robustness and learning theoretic implications on optimal stochastic control under discounted or average cost criteria. (iii) We thus show that this topology possesses several properties making it ideal to study optimality and approximations (under the weak formulation) and robustness (under the strong formulation) for many applications.
@@ 1 Weak and Strong Topologies on Stochastic Kernels
[…]
@@
Our primary area of application will be control theory: In both deterministic and stochastic control theory, such stochastic kernels play a crucial role, representing the dynamics of a stochastic system ([35, Lemma 1.2], [21, Lemma 3.1], or [4, Lemma F]) or the sets of relaxed/randomized control policies ([77],[31, 45, 44],[20]).
@@ 1.1 Summary of Main Results
Figure 1: Hierarchy of topologies on stochastic kernels and relations with the topologies introduced in our paper: The strong kernel mean embedding topology and weak kernel mean embedding topology: (a) The Young-narrow and Borkar topologies are known in the literature, and their equivalence on the set of stochastic kernels is studied in [79]. (b) In this paper, we introduce the weak kernel mean embedding topology and demonstrate its equivalence to the Young-narrow and Borkar topologies (see Theorem 3.10). (c) We also introduce a strong version of the kernel mean embedding topology, which has been used to establish empirical consistency results for approximating the conditional mean embeddings of stochastic kernels [73, 51]. It clearly implies the weak version of the kernel mean embedding topology. (d’) As established in Theorem 4.1, convergence in the weak kernel mean embedding topology does not imply convergence in the strong version. (d) However, as shown in Theorem 4.2, under the assumption of bounded and equicontinuous densities for the probability measures in the stochastic kernels, we can establish the implication from the weak to the strong version of the kernel topology. (e) Using conditional mean embedding framework [70], one can view stochastic kernels as operators between RKHS and $L_{2}$-space [42] with operator norm topology. [51, Theorem 3.6] shows the strong kernel mean embedding topology for the case with $q=2$ dominates this topology. (f) Alternatively, stochastic kernels can be viewed as Hilbert-Schmidt operators between an RKHS and an $L_{2}$-space, where the Hilbert-Schmidt norm topology, which is stronger than operator norm topology, is equivalent to the strong kernel mean embedding topology for the case with $q=2$ [51, Theorem 5.1].
@@ 1.1.1 A Weak Formulation
The Young Narrow Topology. Since Young’s seminal paper [77], a particularly prominent approach has involved studying the topologies on Young measures defined by stochastic kernels. In this framework, stochastic kernels are identified with probability measures defined on a product space, maintaining a fixed marginal in the input space – typically taken to be the Lebesgue measure in the context of optimal deterministic control (see [77, 48],[23, Section 2.1],[75, p. 254],[47], [7, Theorem 2.2]).
Thus, under the Young narrow topology, one associates with $\Gamma$ the set of probability measures induced on the product space ${\mathsf{Y}}\times{\mathsf{U}}$ with a fixed marginal $\mu$ on $\mathsf{Y}$}. On this product space, several weak topologies and their equivalence properties have been studied in the literature; see e.g, [8], [7, Theorem 2.2] and [32, Theorem 3.32].
The generalization to stochastic control problems to encompass more diverse input measures has become a common practice, especially in applications involving partially observed and decentralized stochastic control. Many of the references have adopted this approach to define control topologies. For example, this methodology has been used in studies on existence and approximation results related to optimal stochastic control [31, 45, 44], decentralized stochastic control [78, Theorem 5.2] and [65, Section 4], piecewise deterministic optimal stochastic control [12, 11], economics and game theory [50, 49, 7], mean-field control policies [13], and optimal quantization [80, Definition 2.1].
The $w^{*}$ or Borkar Topology. Another notable topology on stochastic kernels is the one introduced by Borkar for randomized controls [20]. This topology, which is further elaborated upon in [3, Section 2.4] and builds on the work in [18], is formulated as a $w^{*}$-topology on stochastic kernels. In this framework, policies are treated as mappings from states or measurements to the space of signed measures with bounded variation, denoted as ${\mathcal{M}}({\mathsf{U}})$, where probability measures ${\mathcal{P}}({\mathsf{U}})$ form a subset. We also refer the reader to [27, 29] for further references on such a $w^{*}$-formulation on stochastic kernels, in particular when instead of countably additive signed measures, finitely additive such measures are also considered.
@@
Under the Borkar topology, one studies the space $\Gamma$ using a $w^{*}$-topology formulation as a bounded subset of the set of maps from $\mathsf{Y}$} to the space of signed measures with finite variation, which can be viewed as the topological dual of continuous functions that vanish at infinity. Therefore, according to the Banach-Alaoglu theorem [33, Theorem 5.18], this leads to a relatively sequentially compact metric space. Specifically, the unit ball of $L_{\infty}(\mu,{\mathcal{M}}({\mathsf{U}}))=\mathcal{L}_{1}(\mu,C_{0}({\mathsf{U}}))^{*}$ is compact under the $w^{*}$-topology, resulting in a relatively compact metric topology on stochastic kernels. It is important to note that the presentations in [20, Section 3] and [3, Section 2.4] differ slightly, although the induced topology remains identical. An equivalent representation of this topology is provided in [3, Lemma 2.4.1] (see also [20, Lemma 3.1]).
@@
For our weak formulation, we will obtain several structural properties and investigate its connections with the Young narrow topology and Borkar (or $w^{*}$)-topology and establish equivalence properties with both the Young narrow topology and the $w^{*}$-topology on the set of stochastic kernels. We demonstrate that the new topology we introduce is, in fact, equivalent to both the Young narrow topology and the $w^{*}$-topology on the set of stochastic kernels. This equivalence will be rigorously defined later in the paper, but the idea is that while these topologies may appear slightly different in their construction, they share fundamental properties when applied to stochastic kernels. Such a result is significant because each of these topologies offers distinct advantages in various contexts. For example, based on the results we present, both the $w^{*}$-topology and the kernel mean embedding topology are relatively (sequentially) compact, which is a crucial property when proving existence results. However, they are not closed. In contrast, the Young narrow topology is closed but lacks relative (sequential) compactness. Interestingly, to make the $w^{*}$-topology and the kernel mean embedding topology closed, a tightness condition is required – this ensures uniform concentration of probability measures on compact sets. Similarly, Prokhorov’s theorem shows that the same tightness condition is necessary for the Young narrow topology to achieve relative compactness. Thus, in terms of proving existence results via Weierstrass’s extreme value theorem, these three topologies are equally advantageous. However, the $w^{*}$-topology offers a valuable duality structure that can be leveraged for certain analyses. In addition, the kernel mean embedding topology possesses a Hilbert space structure, which can facilitate approximations of stochastic kernels via simulations, as shown in the work [51].
@@ 1.1.2 A Strong Formulation
Initially, stochastic kernels were assumed to act as operators between the RKHSs of the input and output spaces. However, recent findings [42] show that the input RKHS must be replaced by an $L_{2}$-space because the conditional expectation of an output function belongs to this $L_{2}$-space, not the input RKHS. This shift leads to an operator norm topology for stochastic kernels. Notably, the strong kernel mean embedding topology dominates the operator norm topology [51, Theorem 3.6]. Alternatively, stochastic kernels can be viewed as Hilbert-Schmidt operators between an RKHS and an $L_{2}$-space. In this context, the Hilbert-Schmidt topology is equivalent to the strong kernel mean embedding topology [51, Theorem 5.1]. While exact representations of stochastic kernels via covariance and cross-covariance operators are generally infeasible [42], certain conditions enable such representations using pseudo-inverses and adjoints of the operators [43, Theorem 5.8]. When these conditions are not met, approximate representations can be constructed using Tikhonov-Phillips regularization techniques [51, Section 7].
@@ 1.2 Notation and Conventions
For a metric space $\mathsf{E}$}, the Borel $\sigma$-algebra is denoted by $\mathcal{E}$. We let $C_{b}({\mathsf{E}})$, $C_{0}({\mathsf{E}})$, and $C_{c}({\mathsf{E}})$ denote the set of all continuous and bounded functions on $\mathsf{E}$}, the set of all continuous real functions on $\mathsf{E}$} vanishing at infinity, and the set of all continuous real functions on $\mathsf{E}$} with compact support, respectively. We always assume that $\mathsf{E}$} is locally compact when working with $C_{0}({\mathsf{E}})$ and $C_{c}({\mathsf{E}})$ to avoid any trivial cases. We also note that when $\mathsf{E}$} is compact, the spaces $C_{b}({\mathsf{E}})$, $C_{0}({\mathsf{E}})$, and $C_{c}({\mathsf{E}})$ coincide. In this case, we simply write $C({\mathsf{E}})$ to denote any of them. Let ${\mathcal{M}}({\mathsf{E}})$ and ${\mathcal{P}}({\mathsf{E}})$ denote the set of all finite signed measures and probability measures on $\mathsf{E}$}, respectively. A sequence $\{\mu_{n}\}$ of finite signed measures on $\mathsf{E}$} is said to converge with respect to total variation distance (see [39]) to a finite signed measure $\mu$ if $\lim_{n\rightarrow\infty}2\sup_{D\in\mathcal{E}}|\mu_{n}(D)-\mu(D)|=0$. A sequence $\{\mu_{n}\}$ of finite signed measures on $\mathsf{E}$} is said to converge weakly (see [39]) to a finite signed measure $\mu$ if $\int_{{\mathsf{E}}}gd\mu_{n}\rightarrow\int_{{\mathsf{E}}}gd\mu$ for all $g\in C_{b}({\mathsf{E}})$. A sequence $\{\mu_{n}\}$ of finite signed measures on $\mathsf{E}$} is said to converge vaguely (see [39]) to a finite signed measure $\mu$ if $\int_{{\mathsf{E}}}gd\mu_{n}\rightarrow\int_{{\mathsf{E}}}gd\mu$ for all $g\in C_{0}({\mathsf{E}})$. Unless otherwise specified, the term ‘measurable’ will refer to Borel measurability.
@@ 2 Bochner Spaces and Their Duals
Similar to the definition of measurable functions, we start with definition of simple functions. A function $f:{\mathsf{Y}}\rightarrow{\mathsf{B}}$ is said to be simple if there exists $b_{1},\ldots,b_{n}\in{\mathsf{B}}$ and $E_{1},\ldots,E_{n}\in{\mathcal{Y}}$ such that $f(y)=\sum_{i=1}^{n}b_{i}\,1_{E_{i}}(y).$ Define the Bochner integral of $f$ with respect to $\mu$ as $\int_{{\mathsf{Y}}}f(y)\,\mu(dy)\triangleq\sum_{i=1}^{n}b_{i}\,\mu(E_{i}).$ A function $f:{\mathsf{Y}}\rightarrow{\mathsf{B}}$ is said to be strongly measurable, if there exists a sequence $\{f_{n}\}$ of simple functions with $\lim_{n\rightarrow\infty}\|f_{n}(x)-f(x)\|_{{\mathsf{B}}}=0$ $\mu$-almost-everywhere. The strongly measurable function $f$ is $p$-Bochner-integrable [26] if $\int_{{\mathsf{Y}}}\|f(y)\|_{{\mathsf{B}}}^{p}\,\mu(dy)<\infty$. Let $L_{p}\bigl(\mu,{\mathsf{B}}\bigr)$ denote the set of all $p$-Bochner-integrable functions from $({\mathsf{Y}},{\mathcal{Y}},\mu)$ to $\mathsf{B}$} endowed with the norm
@@
We now identify the topological dual of $L_{p}\bigl(\mu,{\mathsf{B}}\bigr)$ which is denoted by $L_{p}\bigl(\mu,{\mathsf{B}}\bigr)^{*}$. We start with the definition of weakly-measurable functions from $\mathsf{Y}$} to $\mathsf{B}$$^{*}$. A function $\gamma:{\mathsf{Y}}\rightarrow{\mathsf{B}}^{*}$ is called weakly-measurable [24, p. 18] if the mapping ${\mathsf{Y}}\ni y\mapsto\langle\gamma(y),b\rangle\in\mathbb{R}$ is ${\mathcal{Y}}/{\mathcal{B}}(\mathbb{R})$-measurable for all $b\in{\mathsf{B}}$. Let $\mathcal{L}\bigl(\mu,{\mathsf{B}}^{*}\bigr)$ denote the set of all weakly-measurable functions. Then, we define the following subsets
$\mathcal{L}_{q}\bigl(\mu,{\mathsf{B}}^{*}\bigr)$ $\triangleq\biggl\{\gamma\in\mathcal{L}\bigl(\mu,{\mathsf{B}}^{*}\bigr):\|\gamma\|_{q}\triangleq\left(\int_{{\mathsf{Y}}}\|\gamma(y)\|_{{\mathsf{B}}^{*}}^{q}\,\mu(dy)\right)^{1/q}<\infty\biggr\},\,\,1\leq q<\infty,$ (1)
@@
[24, Theorem 1.5.5] Let $q$ be the conjugate of $p$; i.e., $1/p+1/q=1$. For any $\gamma\in\mathcal{L}_{q}\bigl(\mu,{\mathsf{B}}^{*}\bigr)$ and $f\in L_{p}\bigl(\mu,{\mathsf{B}}\bigr)$, let
@@
If $\mathsf{B}$$^{*}$ is reflexive, then $L_{p}\bigl(\mu,{\mathsf{B}}\bigr)^{*}$ can be identified with set of $q$-Bochner integrable functions $L_{q}\bigl(\mu,{\mathsf{B}}^{*}\bigr)$ [56, Theorem 4.2.26]. Indeed, in this case, by [56, Corollary 4.2.5], we have
@@
that is weakly measurability is equivalent to strong measurability111[56, Corollary 4.2.5] states that weak measurability is equivalent to strong measurability if $B^{*}$ is separable. But if $\mathsf{B}$$^{*}$ is reflexive, then $\mathsf{B}$} is also reflexive. Hence $({\mathsf{B}}^{*})^{*}={\mathsf{B}}$, and so, weak measurability is equivalent to weak measurability. Since $\mathsf{B}$} is assumed to be separable, $\mathsf{B}$$^{*}$ is also separable. Therefore, weak-measurability is equivalent to strong measurability..
@@
$\mathcal{L}_{q}\bigl(\mu,G\bigr)\triangleq\biggl\{\gamma\in\mathcal{L}_{q}\bigl(\mu,{\mathsf{B}}^{*}\bigr):\gamma(y)\in G\text{ }\mu-\text{almost-everywhere}\biggr\}.$
If $G$ is a closed and bounded subset, then $\mathcal{L}_{q}\bigl(\mu,G\bigr)$ is also closed and bounded in $\mathcal{L}_{q}\bigl(\mu,{\mathsf{B}}^{*}\bigr)$, and so, by Banach-Alaoglu Theorem [33, Theorem 5.18], it is compact with respect to weak-topology. Since $L_{p}\bigl(\mu,{\mathsf{B}}\bigr)$ is separable, by [39, Lemma 1.3.2], $\mathcal{L}_{q}\bigl(\mu,G\bigr)$ is metrizable, and so, is also sequentially compact.